OpenStudy (anonymous):

Tom has 7 more silver dollars than Jimmy. Together they have a total of 71 silver dollars. How many silver dollars does Tom have?

OpenStudy (ja1):

x+(x+7)=71

OpenStudy (anonymous):

T = the number of silver dollars Tom has J = the number of silver dollars Jimmy has. T = J + 7 (Tom has 7 more silver dollars than JImmy) T + J = 71 (Together they have 71 silver dollars) replace T in the second equation with the J + 7 from the first, and solve for J. then, put this number into the first equation (basically add 7) and that will tell you how many Tom has...giving you the answer.